OpenStudy (anonymous):

Help please.. Radio signals travel at a rate of (3x10^8) meters per second. How many seconds would it take for a radio signal to travel from a satellite to the surface of the earth if the satellite is orbiting at a height of (3.6x10^9) meters?

OpenStudy (anonymous):

Hey, can u plz help me @Mehek14

Mehek (mehek14):

scientific notation?

OpenStudy (anonymous):

im not sure they don't say anything

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

3.6 since the exponent is 9, move the decimal point 9 units to the right

OpenStudy (anonymous):

3600000000. ? @Mehek14

Mehek (mehek14):

yup don't need the decimal point at the end. :) GJ

OpenStudy (anonymous):

thanks so what do we do next ?

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

time = distance/speed

ganeshie8 (ganeshie8):

you have distance = (3.6x10^9) meters speed = (3x10^8) meters per second simple take the ratio for time : \[\text{time} = \dfrac{3.6\times 10^9}{3\times 10^8}\]

OpenStudy (anonymous):

we're gonna use the rule "PEDMAS" right ?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the exponent rule : \[\large \dfrac{a^m}{a^n} = a^{m-n}\]

ganeshie8 (ganeshie8):

\[\text{time} = \dfrac{3.6\times 10^9}{3\times 10^8} = \dfrac{3.6\times 10^{9-8}}{3} = \dfrac{3.6\times 10^1}{3} = \dfrac{36}{3} = ?\]

OpenStudy (anonymous):

12 seconds, i got it thank you very much !! @ganeshie8